OpenStudy (anonymous):

How many solutions can be found for the equation 5x + 3(x − 1) = 10x − 2x − 3? None One Two Infinitely many

OpenStudy (anonymous):

First solve the equation: 5x + 3(x - 1) = 10x - 2x - 3 Start with 3(x - 1) what is that?

OpenStudy (anonymous):

idk what x is

OpenStudy (anonymous):

your just multiplying it so it would be 3x-3

OpenStudy (anonymous):

um

OpenStudy (anonymous):

the x is confusing me

OpenStudy (anonymous):

first you have to simplify each side which is 5x+3x-3=10x-2x-3 Then you combine like terms so its 8x-3=8x-3. Since both sides are identical, there are infinite solutions

OpenStudy (anonymous):

Now, you have 5x - 3x - 3 = 10x - 2x - 3

OpenStudy (anonymous):

Put all the same type variables on one side: 5x - 3x - 2x + 10x = -3 - 3

OpenStudy (anonymous):

Now simplify that: 10x = -6 Divide 10/6 = 1 2/3 No solution.